Suppose f(x,y)=x2+y2−6x and D is the closed triangular region with vertices (6,0), (0,6), and (0,−6). Answer the following. Find the absolute maximum of f(x,y) on the region D. Find the absolute minimum of f(x,y) on the region D.
1
Expert's answer
2020-04-06T16:30:30-0400

f(x,y)=x2+y2−6x

A (6,0)

B (0,6)

C (0, -6)


fx= 2x-6, fx= 0 -> x = 3

fy= 2y, fy= 0 -> y = 0


Value in critical point f(3,0) = -9

AB: y = 6-x, 0<=x<=6

BC: y = x-6, 0<=x<=6

AC: x= 0, -6<=x<=6


f(x,6-x) = x2+(6-x)2-6x = 2x2-18x+36

f'(x,6-x) = 4x-18, f'(x,6-x) = 0, -> x = 9/2, y = 3/2

Calculate values at the two end points of AB:

f(0,6) = 36

f(9/2,3/2) = -9/2

f(6,0) = 0


Calculate for vector BC:

f(x,x-6) = x2+(x-6)2-6x = 2x2-18x+36

f'(x,6-x) = 4x-18, f'(x,x-6) = 0, -> x = 9/2, y = 3/2

Calculate values at the two end points of BC:

f(0,-6) = 36

f(9/2,3/2) = -9/2

f(6,0) = 0


Finally solve for the AC:

f(0,y)= y2

So f(0,0) = 0


Absolute minimum: f(3,0) = -9

Absolute maximum: f(0,6)=f(0,-6) = 36